suppose that 100 tyres of a certain brand lasted on the average 21431 miles with a standard deviation of 1295 miles. using a level of significance of 0.05, test the hypothesis mean = 22000 miles against the alternative hypothesis mean < 22000 miles.
1
Expert's answer
2021-10-04T16:16:20-0400

"\bar{x}=21431 \\

s=1295 \\

n=100 \\

H_0: \mu = 22000 \\

H_1: \mu < 22000"

Test-statistic:

"t= \frac{\bar{x} - \mu}{s / \sqrt{n}} \\

t = \frac{21431-22000}{1295 / \sqrt{100}} = -4.39 \\

α=0.05 \\

df=100-1 = 99"

Rejection region:

Reject the H0 if P-value is less than 0.05.

P-value:

The P-value is computed using the excel formula for t-distribution which is =T.DIST(-4.39, 99, TRUE), thus the P-value is computed as "1.42 \times 10^{-5}."

Conclusion: Since the P-value is less than 0.05 hence we reject the null hypothesis and conclude that mean < 22000 miles at 0.05 level of significance.